An inheritance of R72 000 is shared among 3 beneficiaries, Banele, Anele and Cele in the ratio of 1:3:4

respectively. What will be the inheritance for Anele, Banele and Cele (in that order) respectively?


1
Expert's answer
2021-05-04T13:08:48-0400

Ratio of sharing inheritance among Banele, Anele and Cele is 1:3:4

So By Ratio

Inheritance Of Anele= 3/8*(R 72000)=R 27000

Inheritance Of Banele=1/8*(R 72000)=R 9000

Inheritance Of Cele= 4/8*(R 72000)= R 36000